A passenger train can travel 325 mi in the same time a freight train takes to travel 200 mi. If the speed of the passenger train is 25 mi/h faster than the speed of the frieght train, find the speed of each.
I came up with 45 mi/h and I am not sure this is not correct, but I need to verify this. Thanks

Answer by rmromero(383)


A passenger train can travel 325 mi in the same time a freight train takes to travel 200 mi. If the speed of the passenger train is 25 mi/h faster than the speed of the frieght train, find the speed of each.
What is asked in the problem?
Find the speed of each
Given :
passenger train can travel 325 mi same time
a freight train takes to travel 200 mi.
the speed of the passenger train is 25 mi/h faster
than the speed of the freight train
Illustration:
d / r = t
Passenger train 325 x + 25 = 325/(x + 2
Freight train 200 x = 200/ x
Time is the same
325 200
_______ = _____
x + 25 x
Cross multiply
325x = 200 (x + 25)
325x = 200x + 5000
325x - 200x = 5000
125x = 5000
x = 5000/125
x = 40 mi/h ----->Speed of the Freight train
x + 25mi/h = 40 + 25
= 65 mi/h ---->Speed of the Passenger Train
If you check they travel 5 hours to reach each distance with
diferent speed.
d = rt
passenger train
325mi = 65mi/h * 5h
325mi = 325 mi -------> True
Freight train
200 mi = 40mi/h * 5h
200 mi = 200 mi ---------> True